OpenStudy (anonymous):

prove the identity. show all work. cos(x-pi/2)=sin x

hartnn (hartnn):

cos (A+B) = ?

OpenStudy (anonymous):

huh?

hartnn (hartnn):

cos (A+B) =cos A cos B - sin A sin B now put A=x and B=pi/2

OpenStudy (anonymous):

but its cos (x-pi/2) not addition?

hartnn (hartnn):

sorry, B = -pi/2 or u can use cos (A-B) = cos A cos B +sin A sin B

OpenStudy (anonymous):

ohh ok so cos xcospi/2+sin x cospi/2

hartnn (hartnn):

cos x cos pi/2+sin x sin pi/2 now what is cos pi/2 and sin pi/2 = ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

sin =1

hartnn (hartnn):

so plug those values in, what u get ?

OpenStudy (anonymous):

so the rhs cancels ans ur left with 1 *sin x which is equal to sinx

hartnn (hartnn):

thats correct :)

hartnn (hartnn):

cos x cos pi/2+sin x sin pi/2 = cos x (0)+sin x (1) = sin x

OpenStudy (anonymous):

yay thank you only a few more questions and im done

hartnn (hartnn):

great ! bring it on ...